# Incorporating Anharmonicity within Density Functional Theory Calculations

### v1

@@ Abstract
We present a simple and efficient method to incorporate anharmonic effects in the vibrational frequency of molecules within density functional theory (DFT) calculations. This approach is closely related to the traditional vibrational complete interaction (VCI) technique, which uses the harmonic oscillator wavefunctions as the basis. In our implementation, we employ Gaussian-type orbitals (GTOs), with polynomial prefactors, as the basis set to evaluate the anharmonic Hamiltonian. Although these basis functions are non-orthogonal, the matrix elements such as overlap, kinetic energy terms, and position moments can be evaluated analytically. The terms in the Hamiltonian due to the anharmonic potentials are numerically calculated on a Hermite-Quadrature grid. The potentials can be evaluated using any electronic structure method. This framework enables us to accurately calculate the anharmonicity-corrected vibrational frequencies, the fundamental frequencies, and the corrections to bond lengths in diatomic molecules. This method is also generalized to handle coupled anharmonic oscillators, which is essential to model more complex phenomena such as nitrogen tunneling in the umbrella mode of ammonia (NH3) and Fermi resonances in carbon dioxide (CO2).
@@ I Introduction
In this paper, we present a method that can be used to accurately evaluate the vibrational frequencies of molecules beyond the harmonic approximation. In the harmonic approximation, the potential energy surface (PES) along each normal mode is assumed to be purely quadratic. However, real molecular vibrations deviate from this idealized behavior, resulting in anharmonicity. Accounting for anharmonicity not only shifts the computed vibrational frequencies closer to experimentally observed values, but also provides a more physically accurate description of molecular behavior, including energy level spacing, mode coupling, and vibrational intensity patternsHrenar, Werner, and Rauhut (2007); Barone, Biczysko, and Bloino (2014). When incorporating anharmonic effects in electronic structure calculations, the most widely used methods include vibrational perturbation theory (VPT2) Barone (2005); Bloino and Barone (2012), vibrational self-consistent field (VSCF)Bowman (1986); Gerber and Ratner (1988), and vibrational configuration interaction (VCI) approaches Christiansen (2007); Rauhut (2007); Erba et al. (2019). The method presented here is equivalent to the VCI method. We will discuss this further in the theory and implementation section.
@@ II Theory and Implementation
where $M^{\prime}$ is the effective mass of the mode. In the VCI method, the quantum harmonic oscillator wavefunctions including higher excited states are used as basis functions for this Hamiltonian. However, we choose to use the simpler product functions shown in equation as our basis functions. These span the same space as the wave functions of the quantum harmonic oscillator.
@@
$t_{n,m}=\frac{1}{2M^{\prime}}\braket{\psi_{n}|\frac{d^{2}}{dx^{2}}|\psi_{m}}=\frac{1}{2M^{\prime}}\int_{-\infty}^{+\infty}x^{n}e^{-A^{2}x^{2}/2}\frac{d^{2}[x^{m}e^{-A^{2}x^{2}/2}]}{dx^{2}}dx$ (5)
@@
However, Eq. 6 can also be rewritten as the sum of two integrals by using an integration by parts. In each integral, the order of the integrand is reduced by 2. Due to this, a smaller number of Hermite-quadrature points is required for the numerical integrations compared to Eq. 6. The NRLMOL DFT software package Pederson and Jackson (1990); Jackson and Pederson (1990) used for this study has the ability to accurately and efficiently calculate the gradients of energy with respect to atomic positions.
@@ III.2.1 Ammonia (NH3): coupling between the umbrella mode and the symmetric stretch mode
NH3 has a pyramidal geometry in which the nitrogen atom stays above the plane of the three H atoms. The umbrella mode of the molecule is influenced by the nitrogen atom tunneling through the plane of the hydrogen atoms to invert the pyramidDennison and Uhlenbeck (1932). The tunneling effect causes splitting of the vibrational energy levels. We utilized our coupled-anharmonic oscillator approach to simulate this effect, by allowing the umbrella and the symmetric bond stretch modes to couple. First, we found the optimal geometry of NH3 when it is flat. This is known as the transition state geometry for N-tunneling reaction. Then, a harmonic vibrational analysis was performed to find normal modes and corresponding eigenvectors for the flat geometry. Hermite grid points are always symmetric around the reference geometry for a given eigenmode. By selecting the transition state geometry of NH3 as the reference, we can generate Hermite grid points that allow integration of the PES along the umbrella mode, including the region containing the two minima along this mode. Fig. 3 represents the energy surface of the molecule with respect to displacements along the eigenvectors (alternatively the vibron displacement vectors) of the umbrella (x) and symmetric stretch (y) modes. This was done with a 16x16 Hermite grid. At $x=0$ (flat), the energy surface along y is a minimum at y=0, while at y=0, the energy surface along x at $x=0$ is a maximum.
@@
We investigated the number of basis functions required for each mode to obtain converged results. Figure 4 presents the vibrational energy in the ground state when a different number of basis functions (4 to 14) is used per mode. These results confirm that the ground-state energy converges with a comparatively small number of basis functions for the symmetric stretch mode, while the bending mode requires a larger basis set to achieve convergence. Table 2 shows how the ground state energy level and the splitting vary with the number of basis functions per mode. For simplicity, the same number of basis functions were used for both modes. The reported experimental splitting for the ground state energy level due to tunneling is $\sim$0.8 cm-1,Benedict and Plyler (1958). Our result is in good agreement with that.
@@
Table 2: Dependence of ground-state, first two excited-state energies relative to ground state on the number of basis functions per mode.
@@ III.2.2 CO2: Fermi-resonance
Vibration of the CO2 molecule is a classic example where anharmonicity leads to an interesting phenomenon referred to as the Fermi-resonanceFermi (1931). CO2 is a linear molecule with 4 vibrational modes. The first overtone frequency of the bending mode (which is doubly-degenerate) happens to be approximately equal to the frequency of the symmetric stretch mode. Both have the same symmetry. Since the actual vibrational potential energy has anharmonic contributions, they interact with each other and produce mixed and split vibrational states. As a result, the vibrational spectrum of CO2 shows additional Raman active modes. This is known as the Fermi resonance in CO2.
@@
Fig. 7 shows the vibrational DOS and the Raman active modes. As expected from the Fermi resonance, the two modes at 1219 cm-1 and 1334 cm-1 are Raman active. Although these values do not exactly match the experimental observationsRasetti (1929); Dickinson, Dillon, and Rasetti (1929); Herzberg (2013) (1285 cm-1 and 1388 cm-1), the splitting between the two modes obtained from our calculation (115 cm-1) is in close agreement with the experimental value (103 cm-1), indicating that our model captures the essence of the anharmonic coupling responsible for the Fermi-resonance in CO2.
@@
Figure 7: Calculated vibrational density of states and approximate Raman activity of the CO2 molecule when all four modes are allowed to couple.
@@ IV Summary
In this work, we presented a framework for implementing the VCI method based on a Gaussian function basis set. Owing to this choice of basis set, overlap matrix elements, kinetic energy terms in the Hamiltonian, and properties such as moments can be evaluated analytically. Only for the PES integration do we employ an efficient numerical grid derived from the Gauss–Hermite quadrature method. The PES were obtained from DFT total energy calculations using the GGA-PBE xc functional. As an initial application, we tested our approach on one-dimensional cases (diatomic molecules) and found that converged vibrational spectra can be obtained with as few as six basis functions. We also found that the corrections to bond lengths due to anharmonicity in these diatomic molecules were important when their constituents are H atom and alkali or alkaline earth metal atoms. We further demonstrated that the method can be generalized to model coupled anharmonic oscillations in molecules containing more than 2 atoms. We allowed the symmetric stretch and umbrella modes to couple in NH3 molecule. By doing that, we were able to demonstrate energy level splitting due to N-tunneling through the plane of the hydrogen atoms. One limitation of the method was that higher order basis functions for the bending mode were required to get converged energy levels. However, this behavior is expected. We used a flat NH3 geometry as the reference geometry, which is a maximum along bending mode. In the CO2 molecule, we showed that additional modes (splitting) occur only when the two bending modes and the symmetric stretch modes were allowed to couple. In the case where we allowed all four modes to couple, we found that there were two Raman active modes at 1219 cm-1 and 1334 cm-1. This behavior is known as Fermi-resonance. These examples demonstrate that our approach captures important physical phenomena that modifies the vibrational spectrum of molecules due to anharmonicity in the PES.

### v2

# Universal Approach for Determining Multi-Dimensional Anharmonic Vibrations from Electronic Quantum Methods

@@ Abstract
We present a simple and efficient method to incorporate anharmonic effects in the vibrational analyses of molecules within density functional theory (DFT) calculations. This approach is closely related to the traditional vibrational configuration interaction (VCI) technique, which uses the harmonic oscillator wavefunctions as the basis. In our implementation, we employ Gaussian-type orbitals (GTOs), with polynomial prefactors, as the basis set to evaluate the anharmonic Hamiltonian. Although these basis functions are non-orthogonal, the matrix elements such as overlap, kinetic energy terms, and position moments can be evaluated analytically. The terms in the Hamiltonian due to the anharmonic potentials are numerically calculated on a Hermite-Quadrature grid. The potentials can be evaluated using any electronic structure method. This framework enables us to accurately calculate the anharmonicity-corrected vibrational frequencies, the fundamental frequencies, and the corrections to bond lengths in diatomic molecules. This method is also generalized to handle coupled anharmonic oscillators, which is essential to model more complex phenomena such as nitrogen tunneling in the umbrella mode of ammonia (NH3) and Fermi resonances in carbon dioxide (CO2).
@@ I Introduction
In the harmonic approximation, the potential energy surface (PES) along each normal mode is assumed to be purely quadratic. However, real molecular vibrations deviate from this idealized behavior, resulting in anharmonicity. Accounting for anharmonicity not only shifts the computed vibrational frequencies closer to experimentally observed values, but also provides a more physically accurate description of molecular behavior, including energy level spacing, mode coupling, and vibrational intensity patternsTomica_anh; barone2014fully. When incorporating anharmonic effects in electronic structure calculations, the most widely used methods include vibrational perturbation theory (VPT2) barone2005anharmonic; 10.1063/1.3695210, vibrational self-consistent field (VSCF)bowman1986self; gerber1988self, and vibrational configuration interaction (VCI) approaches christiansen2007vibrational; rauhut2007configuration; doi:10.1021/acs.jctc.9b00294. The method presented here is equivalent to the VCI method. We will discuss this further in the theory and implementation section.
@@ II Theory and Implementation
where $M^{\prime}$ is the effective mass of the mode. In the VCI method, the quantum harmonic oscillator wavefunctions including higher excited states are used as basis functions for this Hamiltonian. However, we choose to use the simpler product functions shown in equation 3 as our basis functions. These span the same space as the wave functions of the quantum harmonic oscillator.
@@
$t_{n,m}$ $=\frac{-1}{2M^{\prime}}\braket{\psi_{n}|\frac{d^{2}}{dx^{2}}|\psi_{m}}$
$=\frac{-1}{2M^{\prime}}\int_{-\infty}^{+\infty}x^{n}e^{-A^{2}x^{2}/2}\frac{d^{2}[x^{m}e^{-A^{2}x^{2}/2}]}{dx^{2}}dx$
$={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{-1}{2M^{\prime}}\Big[m(m-1)s_{n,m-2}-A(2m+1)s_{n,m}+A^{2}s_{n,m+2}\Big]}$ (5)
@@
However, Eq. 6 can also be rewritten as the sum of two integrals by using an integration by parts. In each integral, the order of the integrand is reduced by 2. Due to this, a smaller number of Hermite-quadrature points is required for the numerical integrations compared to Eq. 6. The NRLMOL DFT software package pederson1990variational; jackson1990accurate used for this study has the ability to accurately and efficiently calculate the gradients of energy with respect to atomic positions.
@@ III.2.1 Ammonia (NH3): coupling between the umbrella mode and the symmetric stretch mode
NH3 has a pyramidal geometry in which the nitrogen atom stays above the plane of the three H atoms. The umbrella mode of the molecule is influenced by the nitrogen atom tunneling through the plane of the hydrogen atoms to invert the pyramiddennison1932two. The tunneling effect causes splitting of the vibrational energy levels. We utilized our coupled-anharmonic oscillator approach to simulate this effect, by allowing the umbrella and the symmetric bond stretch modes to couple. First, we found the optimal geometry of NH3 when it is flat. This is known as the transition state geometry for N-tunneling reaction. Then, a harmonic vibrational analysis was performed to find normal modes and corresponding eigenvectors for the flat geometry. Hermite grid points are always symmetric around the reference geometry for a given eigenmode. By selecting the transition state geometry of NH3 as the reference, we can generate Hermite grid points that allow integration of the PES along the umbrella mode, including the region containing the two minima along this mode. Fig. 3 represents the energy surface of the molecule with respect to displacements along the eigenvectors (alternatively the vibron displacement vectors) of the umbrella (x) and symmetric stretch (y) modes. This was done with a 16x16 Hermite grid. At $x=0$ (flat), the energy surface along y is a minimum at y=0, while at y=0, the energy surface along x at $x=0$ is a maximum.
@@
We investigated the number of basis functions required for each mode to obtain converged results. Figure 4 presents the vibrational energy in the ground state when a different number of basis functions (4 to 14) is used per mode. These results confirm that the ground-state energy converges with a comparatively small number of basis functions for the symmetric stretch mode, while the bending mode requires a larger basis set to achieve convergence. Table 2 shows how the ground state energy level and the splitting vary with the number of basis functions per mode. For simplicity, the same number of basis functions were used for both modes. The reported experimental splitting for the ground state energy level due to tunneling is $\sim$0.8 cm-1,benedictvibration. Because of the poor convergence in the calculated splitting, we further investigated this by adding more basis functions. Hermite grids of 20x20 and 24x24 were utilized and the PES was calculated on them for the numerical integration. In both cases, basis functions up to $x^{20}e^{-A^{2}x^{2}/2}$ were added in both the umbrella and symmetric stretch modes. Then, we noticed that the calculated splitting converges to $\sim 0.84$ cm-1.
@@
Table 2: Dependence of ground-state, first two excited-state energies relative to ground state on the number of basis functions per mode for NH3
@@ III.2.2 CO2: Fermi-resonance
Vibration of the CO2 molecule is a classic example where anharmonicity leads to an interesting phenomenon referred to as the Fermi-resonancefermi1931ramaneffekt. CO2 is a linear molecule with 4 vibrational modes. The first overtone frequency of the bending mode (which is doubly-degenerate) happens to be approximately equal to the frequency of the symmetric stretch mode. Both have the same symmetry. Since the actual vibrational potential energy has anharmonic contributions, they interact with each other and produce mixed and split vibrational states. As a result, the vibrational spectrum of CO2 shows additional Raman active modes. This is known as the Fermi resonance in CO2.
@@
Fig. 7 shows the vibrational DOS and the Raman active modes. As expected from the Fermi resonance, the two modes at 1219 cm-1 and 1334 cm-1 are Raman active. Although these values do not exactly match the experimental observationsrasetti1929raman; Dickinson1929; herzberg2013molecular (1285 cm-1 and 1388 cm-1), the splitting between the two modes obtained from our calculation (115 cm-1) is in close agreement with the experimental value (103 cm-1), indicating that our model captures the essence of the anharmonic coupling responsible for the Fermi-resonance in CO2.
@@
Refer to caption 667 1285 1388 2349
Figure 7: Calculated vibrational density of states and approximate Raman activity of the CO2 molecule when all four modes are allowed to couple. The experimental peaks are indicated using vertical dashed lines in black.
@@ IV Summary
In this work, we presented a framework for implementing the VCI method based on a Gaussian function basis set. Owing to this choice of basis set, overlap matrix elements, kinetic energy terms in the Hamiltonian, and properties such as moments can be evaluated analytically. Only for the PES integration do we employ an efficient numerical grid derived from the Gauss–Hermite quadrature method. The PES were obtained from DFT total energy calculations using the GGA-PBE xc functional. As an initial application, we tested our approach on one-dimensional cases (diatomic molecules) and found that converged vibrational spectra can be obtained with as few as six basis functions. We also found that the corrections to bond lengths due to anharmonicity in these diatomic molecules were important when their constituents are H atom and alkali or alkaline earth metal atoms. We further demonstrated that the method can be generalized to model coupled anharmonic oscillations in molecules containing more than 2 atoms. We allowed the symmetric stretch and umbrella modes to couple in NH3 molecule. By doing that, we were able to demonstrate energy level splitting due to N-tunneling through the plane of the hydrogen atoms. However due to the complexity of the PES along the umbrella mode, higher order basis functions are required to achieve convergence in both the ground-state energy and and splitting. This requires a denser Hermite grid for accurate PES integration. We view this behavior as a limitation of the method. In the CO2 molecule, we showed that additional modes (splitting) occur only when the two bending modes and the symmetric stretch modes were allowed to couple. In the case where we allowed all four modes to couple, we found that there were two Raman active modes at 1219 cm-1 and 1334 cm-1. This behavior is known as Fermi-resonance. These examples demonstrate that our approach captures important physical phenomena that modifies the vibrational spectrum of molecules due to anharmonicity in the PES.
The cost of a full harmonic vibrational calculation when performed thru a series of SCF calculations is proportional to $T_{harm}=(6N)\times I_{e}\times N^{3}\times B^{3}$$(3N)^{3}$. Here N is the number of atoms and B represents the number of basis functions per atom. The number of iterations required per displacement ($I_{e}$) is approximate 6 since wavefunctions from the previous displacement are an excellent starting point for the next displacement. The cost of a full harmonic vibrational calculation followed by a complete anharmonic analysis, on a limited number ($N_{A}\leq 3N$) of the harmonic modes is given by: $T_{anh}^{full}=T_{harm}+(2\sum_{i=1}^{N_{A}}b^{i}_{A})\times I_{e}\times N^{3}\times B^{3}+[\Pi_{i=1}^{N_{A}}b^{i}_{A}]^{3}$. Here, $b_{i}$ is at least 2 and represents the mesh size needed for the gaussian Hermite grid (which for the case of NH3 turned out to be 24. For this case, it is not the extra number of SCF calculations that is the rate determining step. The rate determining step becomes the last term $\Pi_{i=1}^{N_{A}}b^{i}_{A}]^{3}$ as it is significantly greater than $2^{3N_{A}}$. So, for example if one chose the seemingly small number of modes ($N_{A}=10$), the cost of the vibrational diagonalization using our approach would be $2^{30}$ which is obviously awful scaling (AS). What our approach can allow us to eventually do is find a smaller number of partner modes, by identifying nearly degenerate pairs of higher harmonics, for each of the $N_{A}$ interesting degrees of vibrational freedom (maybe 3-4) and then perform a series of incomplete low-scaling anharmonic diagonalizaton on each of the $N_{A}$ modes and then use a small number of states from each of these calculations to perform a final full-scale anharmonic calculation at a diagonalization cost that scales as $N^{3}$. Many details needed to be attended to to make this happen. However, we note above identification of the limited number ($N_{A}$) of vibrational modes requires excellent chemical intuition or the initial harmonic analysis but the resulting complexity of a NEO-like analysis of a multi-dimensional vibrational space would have the same AS as our complete method albeit with an amplifying prefactor of $I_{e}$. Perhaps there is a similar incomplete method for coupled methods but these details also need to be worked out. Multiple methods are needed to (1) reduce $b_{i}$, for which there are prescriptions, and to automatically and systematically identify the interesting vibrational degrees of freedom.
Anharmonic vibrational effects can be treated at either the Born-Oppenheimer approximation, at the level of configuration interaction, or in more complicated cases the well established nuclear-electronic orbital method in which case both electrons and nuclei are treated quantum mechanically. Dating back to early work by theoretical work by Goddard SURRATT197739 (including references therein) and experimental work by Herzberg Herzberg; Herzber46, there is a history of work on the methyl (CH3) radical. From computational and theoretical perspectives Born-Oppenheimer effects and multi-configurational effects have been examined in attempting to understand the double- versus single- well approximations in this case. An analysis of that series of papers suggests that one can not necessarily criticize a lower-level approach which yields a Born-Oppenheimer double well just because a higher level predicts a single well and instead one must look at the reliability of the predicted vibrational spectra. The reason for this is that total energy, which includes purely electronic and anharmonic vibrational terms in the former case compared to a single total energy in the second case leads to the fact that the resulting energy vs displacement curves do not distinguish between standard zero-point contributions and electronic contributions so the appearance of single-well nature in the total energy does not contradict BO results. This fact has been overlooked in more recent cases that argue for spurious symmetry breaking for a new density functional method for the case of the methyl radical Hahn. Perhaps because of such issues, Hammes-Schiffer and others have argued for explicit product functions that include the nuclear and electronic wavefunction NEOSHS; SHSXLi and there are of course countless other examples where gaussians have been used for multi-center basis functions. At one level one could argue that an explicitly self-consistent product representation is better than Born-Oppenheimer approaches as it adds nuclear quantum effects in a manner that does not significantly change scaling vs system size and only adds the appearance of a diagonalization of the standard electronic Hamiltonians, for both spins, that is performed at the same time as a much smaller solution to the vibrational wavefunctions. However effectively avoiding non-systematic errors due to the electron-nuclear cusp condition CuspRJNeeds is guaranteed in a Born-Oppenheimer based approach which renders a plug-and-play software for to all electronic-structure methods - particularily all-electron software. However, in regard to any method describing vibrational effects of nuclei, it is the case that when one is interested in accounting for full anharmonicity involving many nuclear degress of freedom, the time required for simultaneous diagonalization of the vibrational quickly accelerates O($M^{3}$)O($N^{3}$), with M and N the number of basis functions per mode and atoms respectively. For the general case there are not a priori approaches for conjuring up which modes, to select from the 3N-possible phonon eigenstates, to address more perfectly. The method proposed determines the interesting anharmonic/vibrational space from the initial vibrational calculation porezag. Here we argue that a post-SCF anharmonic calculation which allows for both identification and treatment of chemically interesting anharmonic couplings between nearly degenerate albeit distinct vibronic states could be coupled to many approaches include NEO. We hope that future studies will use the approach described here to stray further away from the use of standard Hermite polynomials and the resulting Gauss-Hermite-Quadrature approaches to a prescription that instead uses the the resulting space of vibrational eigenstates determined for the one-dimensional space to generate reduced grids and reduced basis functions.
Supplementary Material
See the Supplementary Material for the reference geometries, normal-mode eigenvectors, Gauss–Hermite grid points, and corresponding weights for NH3 and CO2 used in the anharmonic analyses.